Using the addition method, what is the solution to 2x-5y=4 and 7x-2y=4
I've tried and I can't figure it out. I eliminate the x's by multiplying to 14 but i get the fraction of 16/31y and it isn't correct.

Answer by stanbon(75887)

Using the addition method, what is the solution to
2x-5y=4
7x-2y=4
--------------------------------------
Multiply thru the 1st equation by 2; Multiply thru the 2nd equation by 5.
---
4x-10y = 8
35x-10y = 20
--------------------
Subtract the 1st from the 2nd to solve for "x":
31x = 12
x = 12/31
------------------------
Substitute into 2x-5y=4 to solve for "y":
2(12/31) - 5y = 4
24/31 - 5y = 124/31
5y = -100/31
y = -20/31
--------------------
Cheers,
Stan H.